A certain quantity of gas, when compressed to a volume of 0.82 m^3 has a pressure of 76Pa. Find the pressure of the gas when it is compressed to 0.55. Assume a constant temperature.
Answer by venugopalramana(3286)

A certain quantity of gas, when compressed to a volume of 0.82 m^3 has a pressure of 76Pa. Find the pressure of the gas when it is compressed to 0.55. Assume a constant temperature.
AS PER BOYLES LAW at constant temperature PRODUCT OF PRESSURE AND VOLUME REMAINS CONSTANT ...SO.....
76*0.82=P*0.55
P=76*0.82/0.55 = 113.3 Pa